OpenStudy (anonymous):

What value of k solves the equation? k^-5 = 1/32

OpenStudy (anonymous):

\[\frac{ 1 }{ k ^{5} } =\frac{ 1 }{ 32 }\]

OpenStudy (anonymous):

32= \[k^5\]

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):

\[\sqrt[5]{32}=k\]

OpenStudy (anonymous):

\[k=2\]

OpenStudy (anonymous):

a medal please